OpenStudy (anonymous):

solve the question: 4(d-3)=2d and s+2-3s-16=0

OpenStudy (anonymous):

do you know how to distribute numbers into parenthesis?

OpenStudy (anonymous):

this is how i solved it before i got lost 4d-12=2d -2 -2 2d - 12

OpenStudy (anonymous):

almost right, you did it in your head correctly but not on paper 4d-12=2d -2d -2d 2d-12=0

OpenStudy (anonymous):

you weren't missing much so far,

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

ohhhh ok ... the second one s+2-3s-16=0 +3 +3 4s+2-16=0 4s-14=0 +14 +14 4s = 14 /4 /4

OpenStudy (anonymous):

s+2-3s-16=0 reverse distribute (factoring) (1-3)*s +2-16=0

OpenStudy (anonymous):

following?

OpenStudy (anonymous):

mhmm so it has to equal 0?

OpenStudy (anonymous):

no, let's do this step-by-step

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

this time let's separate the s's from the numbers (move numbers to other side of equals sign s+2-3s-16=0 -2 -2 on the left side the 2 and -2 cancel eachother out, so we get s+2-2-3s-16=0-2 s-3s-16=-2

OpenStudy (anonymous):

does this approach to this problem make more sense

OpenStudy (anonymous):

s-3s-16=-2 +16 +16 s-3s-16+16=-2+16 s-3s=14

OpenStudy (anonymous):

ohhh it does. ok .

OpenStudy (anonymous):

sub. s by 3s?

OpenStudy (anonymous):

last two steps are to combine the terms that contain s in them (s-3s) and then dividing by the coefficient in from of s

OpenStudy (anonymous):

that is actually meaning one s minus 3 s

OpenStudy (anonymous):

1s-3s=14

OpenStudy (anonymous):

ok so -2s=12 /12 /12 s=-6?

OpenStudy (anonymous):

sooooooo close

OpenStudy (anonymous):

opssss

OpenStudy (anonymous):

-2s=14 s=-7

OpenStudy (anonymous):

great, you found the error, good job!

OpenStudy (anonymous):

-2s=14 /14 /14 s= -7 ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

woww that wasnt that hard...omg thanks you

OpenStudy (anonymous):

we are actually dividing both sides by -2

OpenStudy (anonymous):

good luck